OpenStudy (anonymous):

Solve for x. x^2=log(x+5) x= ?

OpenStudy (anonymous):

So far I have come up with x+5=10^x^2

OpenStudy (anonymous):

but not sure if I am on the right track

OpenStudy (anonymous):

good to see you again roshan

OpenStudy (anonymous):

Thanks

OpenStudy (anonymous):

Let me help u with this problem then

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

Now use log again

OpenStudy (anonymous):

write log on both sides

OpenStudy (anonymous):

Wait wait This problem is really creeping me out

OpenStudy (anonymous):

hehe

OpenStudy (anonymous):

yes giving me issues as well since I am relearning

OpenStudy (anonymous):

Yeap, but I am trying so wait for some time

OpenStudy (anonymous):

no worries

OpenStudy (anonymous):

I am afraid but the only way to solve this equation is by drawing graphs of lox (x+5) and x^2

OpenStudy (anonymous):

hmmm what do you come up with?

OpenStudy (anonymous):

By drawing graph I came up with x=-1.16 and 1.36

OpenStudy (anonymous):

it does say using a graphing calculator so you may be on the right track. However it wants it to the 4th decimal place.

OpenStudy (anonymous):

Let me try as well and see if I match you

OpenStudy (anonymous):

Allrite

OpenStudy (anonymous):

hmmm what are you graphing? I am not coming up with the same numbers

OpenStudy (anonymous):

I am graphing y=x^2 and y=log(x+5) Now search for the points of intersection

OpenStudy (anonymous):

the two points they cross each other correct?

OpenStudy (anonymous):

yeap

OpenStudy (anonymous):

man I am off for sure

OpenStudy (anonymous):

I blame my calculator lol

OpenStudy (anonymous):

Can u see now?

OpenStudy (anonymous):

yes... wondering why I am so far off... and wishing that sight rounded to 4th decimal lol

OpenStudy (anonymous):

Ha Ha............It was really a tough one, I will try to find another way to solve it as well, and will suggest u the answer